Circular No. 6146

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

CH CYGNI T. Iijima, Asiago Astrophysical Observatory, reports: “A prismatic

spectrogram (range 420–510 nm, resolution 0.3 nm at 500 nm) of this symbiotic star, taken at Asiago Observatory on Mar. 1.2 UT, shows that the emission lines of H I and He I (447.1 nm) have P-Cyg profiles, with absorption features blueshifted by ∼ −500 km/s. On a spectrum taken one day earlier (Feb. 28.1), the emission line of He I (447.1 nm) had a weak absorption component, but H I did not show absorption. A new mass ejection may have occurred abruptly. In addition, the highly-blueshifted broad emission components of the H I lines, first reported by Kuczawska et al. (1992, IBVS 3806) and Iijima et al. (1994, A.Ap. 283, 919), have reappeared. The radial velocities of the broad component of Hβ are −1420 and −880 km/s (double peak) on Feb. 28.1 and −1240 km/s on Mar. 1.2.”

NOVA AQUILAE 1995 J. Johnson, University of Wisconsin, and the Wisconsin Ultraviolet

Photopolarimeter Experiment team (part of the Astro-2 mission aboard the space shuttle Endeavour) report an observation of N Aql 1995: “We obtained a spectrogram (range 140–320 nm, resolution ∼ 1.6 nm) on Mar. 4.84 UT. The spectrum is dominated by Mg II at 280.0 nm. N III] at 175.0 nm, the C IV doublet at 154.8 and 155.0 nm, and the C III] doublet at 190.7 and 190.9 nm are also strong. A line at 164.0 nm could be either He II A at 164.0 nm or O I at 164.1 nm. The O III Bowen fluorescence line at 313.3 nm is seen, and perhaps the lines at 283.9 and 304.7 nm as well. The continuum rises toward longer wavelengths after a possible minimum near 220.0 nm. Another observation was planned for near Mar. 6.33 in order to obtain ultraviolet spectropolarimetry.”

Visual magnitude estimates by P. Schmeer, Bischmisheim, Germany: Feb. 21.206 UT, 9.5; 22.196, 9.4; 27.157, 9.7; Mar. 4.146, 10.0; 5.19, 10.0.

SUPERNOVA 1995H IN NGC 3526 R. H. McNaught, Anglo-Australian Observatory, reports the following

accurate position for SN 1995H: α = 11^{h}06^{m}55^{s}.42, δ = +7^{o}10′19′′.6 (equinox 2000.0; uncertainty 0′′.4 in each coordinate); the corresponding offsets from the galaxy’s center are 19′′.7 west and 7′′.4 south.

1995 March 7 Daniel W. E. Green